OpenStudy (anonymous):

can you use the mean of a x,y table to find the slope and y-intercept or do you have to use rise/run?

OpenStudy (anonymous):

yes you can do it

OpenStudy (anonymous):

tell me the x y numbers and i will help you :)

OpenStudy (anonymous):

Find the equation of the line of best fit from the data in the table. Round the slope and the y-intercept to the nearest hundredth

OpenStudy (anonymous):

i did 12+14+16+18+20=80 80/5=16 54+53+55+54+56=272 272/5=54.4 y=mx+b y=16x+54

OpenStudy (ranga):

Are you sure you are not asked to find the regression line or "the least square fit" line for this data?

OpenStudy (anonymous):